# Targeted Immunisation Thresholds for the Contact Process on Power-Law Trees

### v1

@@ 2.2 Branching random walk and supermartingale
A well-known upper-bound for the contact process is obtained by replacing every infection event by the creation of a new “particle of infection” that further evolves independently as every other particles. We then obtain a simple branching random walk upper bounding the contact process, and in which several “particles of infection” can independently coexist on a single vertex. Using this upper bound, we can easily show that:
  • On a network with degrees strictly less than $1/\lambda$, the total number of infected particles is a subcritical branching process, whose expectation decays exponentially.
  • On a tree with degrees strictly less than $1+1/4\lambda^{2}$, the associated branching process does not survive locally, and thus there is no local survival of the contact process.
@@
One interpretation of the fact that $M^{y}$ is a supermartingale is that starting from a given infected vertex $x$, you only have probability $2\lambda$ to succeed in infecting each vertex on the path from $x$ to $y$, which we could relate to the statement $T(\lambda,1/8\lambda^{2})\leq 2$. In the regime $\tau<2.5$ where $T(\lambda,\mathbbm{t})\to 1$, we will require a straightforward generalization of this supermartingale.
@@
On a tree $\mathcal{T}$ with degrees bounded by $\delta/\lambda^{2}$ with $\delta\leq 1/4$, the processes
@@
Note that $2/(1+\sqrt{1-4\delta})$ tends to 1 as $\delta\to 0$, which is in line with the statement $T(\lambda,\mathbbm{t})\to 1$ if $\mathbbm{t}=o(1/\lambda^{2})$. One may also compare this upper bound with the trivial lower bound
@@
to see that the BRW and CP hitting probabilities are approximately the same for small $\delta$.
These techniques are simple and useful when the degrees are bounded by $1/4\lambda^{2}$ but break down if there is a single vertex of degree $k>1/\lambda^{2}$, as the branching random walk will then be supercritical and thus possibly survive forever, which is not in line with the fact that $T(\lambda,k)=e^{\Theta(\lambda^{2}k)}$ is then possibly large, but still finite.
@@ 2.3 Contact process–Branching random walk
In order to tackle the case of larger degrees, we use the following simple refinement of the branching random walk upper bound, inspired by [periodic2020], where the branching phenomenon is considered only on a specific subset of all the vertices.
Definition 2.2 (CP-BRW with sinks $\mathcal{V}$).
Select a set of “sinks” $\mathcal{V}\subseteq V\left(\mathcal{G}\right)$. Starting from an infection at some sink $v\in\mathcal{V}$, run the contact process, except that every time you observe the infection event of some other sink $w\in\mathcal{V}\backslash\{v\}$, you replace this infection by the creation in $w$ of a “new type of infection” at $w$, that evolves as a contact process but independently of all the previous infections. We then say $v$ “sends an infection” to $w$; note that $v$ may send several infections to $w$ as well as to the other sinks, but none to itself. Doing so, you obtain an upper bound for the contact process, which we call “contact process-branching random walk”, or CP-BRW. In the following, instead of fully studying this continuous-time process and introducing appropriate notations for this purpose, we will lose track of the time, and just count the number of infections that a sink $v$ will eventually send to other sinks $w\in\mathcal{V}\backslash\{v\},$ then the infections that $w$ will send to the sinks in $\mathcal{V}\backslash\{w\}$, etc. In doing so, we obtain a discrete-time branching random walk, which we will denote by $\left(\phi^{\mathcal{V}}_{n}\right)_{n\in\mathbb{N}}$ taking values $\phi^{\mathcal{V}}_{n}:\mathcal{V}\rightarrow\mathbb{N}$.
@@
  • Alternatively, one may imagine that all the infection events that an infection sends to other sinks are “frozen”. Running one step of the BRW construction then consists of independently running a contact process from each of the frozen infections and freezing another for the next step when each other sink is infected.
  • For $\mathcal{V}=V(\mathcal{G})$, the CP-BRW is no different from the (continuous-time) branching random walk mentioned earlier. More generally, adding sinks allows to replace the dynamics of the contact process with the easier dynamics of a branching process, however with the possible drawback of obtaining too crude upper bounds if we introduce too many sinks.
We will often want to bound the probability of the infection hitting a given vertex $v$. To this end, we first define $\xi^{[v]}$ the contact process with absorption at $v$, defined similarly to the contact process except that the infections are absorbed at $v$, and $v$ counts the total number of received infections. Thus $\xi^{[v]}_{t}(v)$ is the total number of infections received at $v$ by time $t$, and
$\{\exists t,\xi_{t}(v)=1\}=\{\xi^{[v]}_{\infty}(v)\geq 1\}.$
To bound the probability of this event, we introduce the CP-BRW with target sink $v$, defined similarly as the CP-BRW, but again with absorption at $v$.
Definition 2.4.
The CP-BRW with sinks $\mathcal{V}$ and target sink $v\in\mathcal{V}$, denoted
$\phi^{\mathcal{V},[v]}:\mathcal{V}\rightarrow\mathbb{N}$
is the CP-BRW of definition 2.2 with the modification that walkers are absorbed at $v$. That is to say, while other sinks will never put offspring on themselves, each walker at $v$ has the deterministic offspring of a single walker at $v$.
Note also that when we have one non-target sink $o$ and one target sink $v$ the CP-BRW has no splitting into types, and thus we have
$\xi_{\infty}^{[v]}(v)=\phi_{1}^{\{o,v\},[v]}(v)=\phi_{\infty}^{\{o,v\},[v]}(v)$
The fact that the CP-BRW provides an upper bound for the contact process can be seen in this context by the following proposition, which follows easily from the graphical construction of the contact process. See also [liggett1985interacting, III.6] for more on the graphical construction and its consequences.
Proposition 2.5.
For $v\in\mathcal{V}\subset\mathcal{W}$ we have a coupling of $\left(\phi_{n}^{\mathcal{V},[v]}\right)$ and $\left(\phi_{n}^{\mathcal{W},[v]}\right)$ such that:
$\phi_{\infty}^{\mathcal{V},[v]}\leq\phi_{\infty}^{\mathcal{W},[v]}.$
When computing the total number of infections absorbed at the target sink $v$ in the CP-BRW, we simply can use the branching structure to obtain, for $o\neq v\in\mathcal{V}$:
$\mathbb{E}_{\mathbbm{1}_{o}}\left(\phi_{\infty}^{\mathcal{V},[v]}(v)\right)=\sum_{\begin{subarray}{c}x_{0}=o,\ x_{k}=v\\ x_{1}\dots x_{k-1}\in\mathcal{V}\setminus\{v\}\\ \forall i,\ x_{i}\neq x_{i-1}\end{subarray}}\prod_{i=1}^{k}\mathbb{E}_{\mathbbm{1}_{x_{i-1}}}\left(\phi_{1}^{\mathcal{V}}(x_{i})\right).$
To enlighten notations, when the starting infected vertex $o$ is clear, we write $m_{\infty}^{[v]}=\operatorname{\mathbb{E}}_{\mathbbm{1}_{o}}[\xi_{\infty}^{[v]}(v)]$ as well as $m_{\infty}^{\mathcal{V},[v]}=\operatorname{\mathbb{E}}_{\mathbbm{1}_{o}}[\phi_{\infty}^{\mathcal{V},[v]}(v)]$ and $m^{\mathcal{V}}(x,y)=\mathbb{E}_{\mathbbm{1}_{x}}\left(\phi_{1}^{\mathcal{V}}(y)\right),$ so that the CP-BRW provides the upper bound
@@
$m_{\infty}^{[v]}\leq m_{\infty}^{\mathcal{V},[v]}=\sum_{\begin{subarray}{c}x_{0}=o,\ x_{k}=v\\ x_{1}\dots x_{k-1}\in\mathcal{V}\setminus\{v\}\\ \forall i,\ x_{i}\neq x_{i-1}\end{subarray}}\prod_{i=1}^{k}m^{\mathcal{V}}(x_{i-1},x_{i}).$ (3)
@@ 3 General upper bounds on the survival probability
$P_{\rm surv}^{(p,q)}(\lambda)=\begin{cases}O\left(\lambda^{\frac{1}{3-\tau}}\right)\quad\text{when }\tau\leq 2.5,\\ O\left(\frac{\lambda^{2\tau-3}}{\log^{\tau-2}\left(\frac{1}{\lambda}\right)}\right)\quad\text{when }\tau\in(2.5,3],\\ O\left(\frac{\lambda^{2\tau-3}}{\log^{2\tau-4}\left(\frac{1}{\lambda}\right)}\right)\quad\text{when }\tau>3.\end{cases}$ (4)
Writing $P_{\rm surv}^{(q)}(\lambda)=P_{\rm surv}(\lambda)$ the survival probability on the GW tree with offspring $q$, we also have
@@
as the contact process has probability $\lambda/(1+\lambda)$ to infect a given neighbour before the recovery of the root, and starting from that point it has probability at least $P_{\rm surv}^{(q)}(\lambda)$ to survive in the subtree containing this neighbour (see also the inequality just after (4.11) in [mountford2013metastable] for a similar statement). The result follows immediately.
@@ 4.1 Supercritical truncation threshold
For $\tau\in(2,2.5]$ this branching process is simply the classical SIR lower bound. For the SIR process, if a vertex with offspring $D\sim\tilde{q}$ ever gets infected, it stays infectious for a period $I\sim\operatorname{Exp}(1)$ and then is removed. During the infectious period it infects a random number $Y$ of its children, which has conditional distribution
$Y\big{|}D,I\sim\operatorname{Bin}\left(D,1-e^{-\lambda I}\right).$
The vertices ever infected form a Galton-Watson tree with offspring distribution $Y$, with mean
@@
For $\tau\in(2,2.5]$, define the critical value $\rho_{c}$ as $c_{q}\frac{\tau-2}{3-\tau}\rho_{c}^{3-\tau}=1$ or equivalently
$\rho_{c}=\left(\frac{3-\tau}{(\tau-2)c_{q}}\right)^{\frac{1}{(3-\tau)}},$ (6)
@@
By a slight refinement of our proof that we do not detail, we also get that the contact process typically survives with probability $\Theta(1)$ when the root is a big vertex with degree $\Theta\left(\lambda^{-\frac{1}{3-\tau}}\right)$, an event which has probability $\Theta\left(\lambda^{\frac{\tau-2}{3-\tau}}\right)$.
@@
$\operatorname{\mathbb{E}}[D]=\sum_{k=1}^{\mathbbm{t}}\operatorname{\mathbb{P}}(D\geq k)\sim\int_{1}^{\mathbbm{t}(\lambda)}c_{q}(x^{-(\tau-2)}-\mathbbm{t}^{-(\tau-2)})\mathrm{d}x\sim\lambda^{-1}c_{q}\frac{\tau-2}{3-\tau}\rho^{3-\tau},$ (7)
and thus by (5) the mean offspring of the Galton-Watson process has $\operatorname{\mathbb{E}}[Y]\to m:=c_{q}\frac{\tau-2}{3-\tau}\rho^{3-\tau},$ which is larger than 1 by the assumption $\rho>\rho_{c}$, and thus the process is supercritical. We further lower bound the survival probability by a simple second moment computation with the following classical lemma.

### v2

@@ 2.3 Branching random walk and supermartingale
It is classical to interpret the contact process as a particle system, and say there is a particle at site $x$ if the site/individual $x$ is infected. In this interpretation, particles are killed at rate $1$ (by recovery of the individual), and give birth with rate $\lambda$ to a new particle at each neighbouring site… with the constraint that the number of particles at each site is bounded by 1 (so giving birth to a particle at an already occupied site has actually no effect). If we remove this constraint and allow multiple particles to coexist at each site, we then obtain a simple branching random walk upper bounding the contact process. This point-of-view is for example explored in detail in [liggett1999stochastic, Part I.4]. The advantage of this upper bound is of course that the branching random walk is often easier to study as the particles now evolve independently. In particular,
  • On a network with degrees strictly less than $\frac{1}{\lambda}$, the total number of infected particles is a subcritical branching process, whose expectation decays exponentially. On the contrary, if the degrees are strictly larger than $\frac{1}{\lambda}$, then the branching random walk is supercritical and has a positive probability of surviving forever (possibly only weakly/globally).
  • On the star graph $S_{k}$ with one central vertex and $k$ leaves, the branching random walk is subcritical if $k<1/\lambda^{2}$ but survives (with positive probability) if $k>1/\lambda^{2}$.
  • On a tree with degrees strictly less than $1+\frac{1}{4\lambda^{2}}$, the branching random walk upper bound does not survive locally [liggett1999stochastic, Theorem 4.8(d)], and thus there is no local survival of the contact process (that is, after some finite time the root will stay healthy forever). On the contrary, on a tree with all degrees strictly larger than $1+\frac{1}{4\lambda^{2}}$, there is strong/local survival.
@@
One interpretation of the fact that $M^{y}$ is a supermartingale is that starting from a given infected vertex $x$, you only have probability $2\lambda$ to succeed in infecting each vertex on the path from $x$ to $y$, which we could relate to the statement $T(\lambda,\nicefrac{{1}}{{8\lambda^{2}}})\leq 2$. In the regime $\tau<2.5$ where $T(\lambda,\mathbbm{t}+1)\to 1$, we will require a straightforward generalization of this supermartingale.
@@
On a tree $\mathcal{T}$ with degrees bounded by $\nicefrac{{\delta}}{{\lambda^{2}}}$ with $\delta\leq\nicefrac{{1}}{{4}}$, the processes
@@
Proof.
Each infected vertex $x$ recovers at rate $1$, and so decreases $M_{t}^{y}$ at rate equal to its weight
$\left(\frac{2\lambda}{1+\sqrt{1-4\delta}}\right)^{d(x,y)}.$
Relative to this weight and by infection, it increases $M_{t}^{y}$ at rate bounded by
$\lambda\cdot\left(\frac{\delta}{\lambda^{2}}-1\right)\cdot\frac{2\lambda}{1+\sqrt{1-4\delta}}<\frac{2\delta}{1+\sqrt{1-4\delta}}$
in its $\left(\nicefrac{{\delta}}{{\lambda^{2}}}-1\right)$ children, and at rate bounded by
$\lambda\cdot\frac{1+\sqrt{1-4\delta}}{2\lambda}=\frac{1+\sqrt{1-4\delta}}{2}$
in its parent. Because
$\frac{2\delta}{1+\sqrt{1-4\delta}}+\frac{1+\sqrt{1-4\delta}}{2}=1$
this shows the supermartingale property. ∎
Note that $2/(1+\sqrt{1-4\delta})$ tends to 1 as $\delta\to 0$, which is in line with the statement $T(\lambda,\mathbbm{t}+1)\to 1$ if $\mathbbm{t}=o(\nicefrac{{1}}{{\lambda^{2}}})$. One may also compare this upper bound with the trivial lower bound
@@
to see that the BRW (branching random walk) and CP (contact process) hitting probabilities are approximately the same for small $\delta$.
These techniques are simple and useful when the degrees are bounded by $\nicefrac{{1}}{{4\lambda^{2}}}$ but cannot treat accurately vertices of larger degrees, as the branching random walk is supercritical if there is a single vertex of degree $k>\nicefrac{{1}}{{\lambda^{2}}}$, while we can still expect that the contact process will not survive around this vertex for time larger than $T(\lambda,k)=e^{\Theta(\lambda^{2}k)}$.
To treat the vertices with higher degrees, we start with simple trees before going to more complicated ones:
  • On a star graph with $k$ leaves, the expected extinction time of the contact process is bounded by $e^{\Theta(\lambda^{2}k)}$.
  • On a tree with one big vertex of degree $k$ and other vertices of degrees bounded by $1/8\lambda^{2}$, we can still provide an upper bound $e^{\Theta(\lambda^{2}k)}$ for the total time that the center is infected, and deduce an upper bound for the hitting probabilities of the vertices in the tree. See Lemma 5.1 and Corollary 5.2.
  • More complicated trees require more work, and in particular the introducing of another process upper bounding the contact process: the “contact process–branching random walk” that we now present.
@@ 2.4 Contact process–Branching random walk
In order to tackle the case of larger degrees, we use the following simple refinement of the branching random walk upper bound, inspired by [periodic2020], where the branching phenomenon is considered only on a specific subset of all the vertices. This upper bound is an intermediate process between the contact process and the branching random walk and shares some features with both, so we will denote it by “contact process-branching random walk”, or CP-BRW. This process is best described as a “multitype contact-process”, with the interpretation of the contact process as an interacting particle system as in the previous section.
Definition 2.3 (continuous-time CP-BRW with sinks $\mathcal{V}$).
Select a set of “sinks” $\mathcal{V}\subseteq V\left(\mathcal{G}\right)$. The continuous-time CP-BRW with sinks $\mathcal{V}$ is a multitype contact process, where each type of infection:
  • is initiated at some sink $v\in\mathcal{V}$.
  • evolves as the usual contact process on $\left(V\left(\mathcal{G}\right)\setminus V\right)\cup\{v\}$. Note this CP has thus at most one particle at each site, and can infect the sink $v$, but not the other sinks.
The different types of particles evolve independently, except for the following rule for the creation of new types of infections:
  • For any sink $v\in\mathcal{V}$, create a new type of infection, initiated at $v$ (and thus starting from only $v$ infected), at a rate given by $\lambda$ times the total number of particles present at the neighbouring sites $z\sim v$, and which have types initiated at other sinks $w\neq v$.
@@
In this description, we stress that the type of an infection is important for two reasons. First, if the type is initiated at $v$, then the contact process evolves on the non-sink sites and also the site $v$. But second and more importantly, its evolution is independent from all the previously created types, even if initiated at the same sites. We further stress that:
  • We typically start the process from one particle, initiated at some sink $v\in\mathcal{V}$.
  • A contact process of given type and initiated at a sink $v\in\mathcal{V}$ cannot imply any creation of new types of infections initiated at $v$. However, it can imply the creation of more than one type of infections initiated at each sink $w\in\mathcal{V}\setminus\{v\}$.
  • To each creation of a new type of infection (say at sink $v$), we can identify the particle which led to this creation (necessarily present at a neighbouring site $z\sim v$), and its type. Referring to this type as “the parent type”, we see that the different types of infections have a natural tree structure, rooted at the original type (the type of the first particle present at time 0).
  • It is not difficult to see that the total number of particles present at any given site (disregarding their types) is a stochastic upper bound for the original contact process starting from only $v$ infected. This is best proven using the graphical construction of the contact process, as constructed and applied in [liggett1985interacting, III.6]. Actually, any increase in the set of sinks means a further stochastic upper bound, and if we consider all sites to be sinks, we just obtain (a complicated description of) the continuous-time Branching random walk of Section 2.3. Heuristically, adding sinks allows to replace the dynamics of the contact process with the easier dynamics of a branching process, however with the possible drawback of obtaining too crude upper bounds if we introduce too many sinks.
Definition 2.5 (discrete-time CP-BRW with sinks $\mathcal{V}$).
For all sinks $w\in\mathcal{V}$ and $n\geq 0$, write $\phi^{\mathcal{V}}_{n}(w)$ for the number of types of infections initiated at $w$ and with depth $n$ in this tree. Then $\left(\phi^{\mathcal{V}}_{n}\right)_{n\in\mathbb{N}}$ is a discrete-time branching-random walk on $\mathcal{V}$, which we will denote as the discrete-time CP-BRW with sinks $\mathcal{V}$, or simply CP-BRW with sinks $\mathcal{V}$.
The discrete-time CP-BRW $\left(\phi^{\mathcal{V}}_{n}\right)_{n\in\mathbb{N}}$ contains much less information than the continuous-time CP-BRW – in particular the time-dynamics are totally lost – but it will be sufficient for our purposes. From now on, if a type of infection initiated at some sink $v\in\mathcal{V}$ has created $k$ types of infections initiated at some other sink $w\in\mathcal{V}$ (namely their children in the tree of the types of the infections), we will simplify the terminology, and simply say that the infection initated at $v$ has “sent $k$ infections” to $w$. We will also write $m^{\mathcal{V}}(v,w)$ for the mean number of infections sent to $w$ by a single infection at $v$, where $v$ need not necessarily be a sink, or simply $m(v,w)$ for $m^{\{w\}}(v,w)$ when there is no other sink.
We will often want to bound the probability of the infection hitting a target sink $w$. To this end, it is natural to consider the CP-BRW with absorption at $w$. It is defined as the CP-BRW in Definition 2.5, except that each walker at $w$ now has the deterministic offpsring of a single walker at $w$. We will denote this process by $\phi^{\mathcal{V},[w]}$, and write $m_{\infty}^{\mathcal{V}}(v,w):=\operatorname{\mathbb{E}}_{\mathbbm{1}_{v}}\left[\phi_{\infty}^{\mathcal{V},[w]}(w)\right]$ for the mean number of infections absorbed at the target sink $w$ when the process is started from a single infection at $v$. Note we can bound the probability that the contact process ever hits $w$ by $m(v,w)\leq m_{\infty}^{\mathcal{V}}(v,w)$. The branching structure of the CP-BRW, allows to express this upper bound as:
@@
$m_{\infty}^{\mathcal{V}}(v,w)=\sum_{\begin{subarray}{c}x_{0}=v,\ x_{k}=w\\ x_{1}\dots x_{k-1}\in\mathcal{V}\setminus\{w\}\\ \forall i,\ x_{i}\neq x_{i-1}\end{subarray}}\prod_{i=1}^{k}m^{\mathcal{V}}(x_{i-1},x_{i}).$ (4)
@@ 3 General upper bounds on the survival probability
$P_{\rm surv}^{(p,q)}(\lambda)=\begin{cases}O\left(\lambda^{\frac{1}{3-\tau}}\right)\quad\text{when }\tau\leq 2.5,\\ O\left(\frac{\lambda^{2\tau-3}}{\log^{\tau-2}\left(\nicefrac{{1}}{{\lambda}}\right)}\right)\quad\text{when }\tau\in(2.5,3],\\ O\left(\frac{\lambda^{2\tau-3}}{\log^{2\tau-4}\left(\nicefrac{{1}}{{\lambda}}\right)}\right)\quad\text{when }\tau>3.\end{cases}$ (5)
Writing $P_{\rm surv}^{(q)}(\lambda)=P_{\rm surv}(\lambda)$ for the survival probability of the contact process on the GW tree with offspring $q$, we also have
@@
as the contact process has probability $\nicefrac{{\lambda}}{{(1+\lambda)}}$ to infect a given neighbour before the recovery of the root, and starting from that point it has probability at least $P_{\rm surv}^{(q)}(\lambda)$ to survive in the subtree containing this neighbour (see also the inequality just after (4.11) in [mountford2013metastable] for a similar statement). The result follows immediately.
@@
In this section we address the parameter regime $\tau\leq 2.5$ of our main theorem by Propositions 4.1 and 4.4.
@@ 4.1 Supercritical truncation threshold
For $\tau\in(2,2.5]$ it suffices to use the classical SIR lower bound of Section 2.1. In a Galton-Watson tree with offspring distribution $D$, the vertices ever infected form a Galton-Watson tree with offspring distribution $Y$, with mean
@@
Recalling from the Introduction that the underlying offspring distribution has $q([k,+\infty))\sim c_{X}k^{2-\tau}$, we have the following result.
@@
For $\tau\in(2,2.5]$, define the critical value $\rho_{c}$ as $c_{X}\frac{\tau-2}{3-\tau}\rho_{c}^{3-\tau}=1$ or equivalently
$\rho_{c}=\left(\frac{3-\tau}{(\tau-2)c_{X}}\right)^{\frac{1}{3-\tau}},$ (7)
@@
By a slight refinement of our proof that we do not detail, we also get that the contact process survives with probability $\Theta(1)$ when the root is a big vertex with degree $\Theta\left(\lambda^{-\frac{1}{3-\tau}}\right)$, an event which has probability $\Theta\left(\lambda^{\frac{\tau-2}{3-\tau}}\right)$.
@@
$\operatorname{\mathbb{E}}[D]=\sum_{k=1}^{\mathbbm{t}}\operatorname{\mathbb{P}}(D\geq k)\sim\int_{1}^{\mathbbm{t}(\lambda)}c_{X}(x^{-(\tau-2)}-\mathbbm{t}^{-(\tau-2)})\mathrm{d}x\sim\lambda^{-1}c_{X}\frac{\tau-2}{3-\tau}\rho^{3-\tau},$ (8)
and thus by (6) the mean offspring of the Galton-Watson process has $\operatorname{\mathbb{E}}[Y]\to m:=c_{X}\frac{\tau-2}{3-\tau}\rho^{3-\tau},$ which is larger than 1 by the assumption $\rho>\rho_{c}$, and thus the process is supercritical. We further lower bound the survival probability by a simple second moment computation with the following classical lemma.